# WiFo-M2: Plug-and-Play Multi-Modal Sensing via Foundation Model to Empower Wireless Communications

### v1

@@ Abstract
The growing adoption of sensor-rich intelligent systems has boosted the use of multi-modal sensing to improve wireless communications. However, traditional methods require extensive manual design of data preprocessing, network architecture, and task-specific fine-tuning, which limits both development scalability and real-world deployment. To address this, we propose WiFo-M2, a foundation model that can be easily plugged into existing deep learning-based transceivers for universal performance gains. To extract generalizable out-of-band (OOB) channel features from multi-modal sensing, we introduce ContraSoM, a contrastive pre-training strategy. Once pre-trained, WiFo-M2 infers future OOB channel features from historical sensor data and strengthens feature robustness via modality-specific data augmentation. Experiments show that WiFo-M2 improves performance across multiple transceiver designs and demonstrates strong generalization to unseen scenarios.
@@ I Introduction
The proliferation of multi-modal sensing, encompassing visual, Light Detection and Ranging (LiDAR), and other forms of environmental data, has become a cornerstone of emerging intelligent systems, from autonomous vehicles and smart factories to pervasive Internet of Things deployments [1]. Correspondingly, the wireless communications community is actively exploring how these ubiquitous sensing sources can be leveraged to optimize wireless communication performance. To systematically give the design principle and underlying mechanisms of sensing-aided communications, Cheng et al. proposed the first framework of this field, Synesthesia of Machines (SoM). To date, numerous studies have investigated promising applications, such as using multi-modal sensors to reduce pilot overhead for channel estimation [3, 4], employing images for beam selection [5, 6, 7], and exploiting on-vehicle sensors to aid multi-user precoding [8, 9, 10].
However, two fundamental challenges hinder the wider adoption and effectiveness of multi-modal sensing in communication systems. First, the prevailing design paradigm remains largely task-specific and scenario-bound. As shown in Fig. 1, Existing approaches typically rely on meticulously customized data processing pipelines and network architectures tailored for narrow objectives, resulting in performance gains that lack universality and entail significant engineering overhead. Second, current research has concentrated primarily on a limited set of transceiver modules (e.g., beam prediction and handover) whose outputs are strongly correlated with visual object detection. The potential of leveraging sensing to enhance a wider range of core transceiver modules, such as channel interpolation and prediction, remains largely untapped due to the complex and often latent relationships between sensing data and various transceiver modules.
To overcome these limitations, a generic method is urgently needed to exploit multi-modal sensing for stable and universal performance gains without extra engineering efforts. This imperative naturally raises the question: How can multi-modal sensing be transformed into task-agnostic feature representations that benefit diverse transceiver modules? Inspired by the emerging paradigm of foundation models (FM) [11, 12], we aim to build one that extracts universal, task-agnostic features representing out-of-band (OOB) channel information. It would serve as a plug-and-play performance booster, adaptable to various transceiver modules.
Figure 1: Comparison between conventional sensing-optimized communications paradigm and the proposed WiFo-M2 paradigm.
@@
To achieve this, we propose WiFo-M2, the first universal foundation model for wireless communication performance enhancement, as shown in Fig. 1. To establish one-to-one positive sample pairs between CSI and multi-modal sensory data, we first develop a unified multi-modal sensing processing pipeline. Then, we design the ContraSoM pre-training strategy, which is designed to align the high-dimensional representations from WiFo-M2 with their corresponding CSI features in the latent space. This alignment allows the WiFo-M2 to extract universally beneficial supplementary features for diverse transceiver modules. To bridge the sampling rate gap, WiFo-M2 is designed as a temporal sequence-to-sequence model. It ingests a fixed historical window of sensory data and is trained to extrapolate future CSI-aligned features. This enables WiFo-M2 to provide supplementary CSI features even when fresh sensory data is temporarily unavailable, a practical issue largely overlooked in prior work. To enhance the robustness and representation learning capability of WiFo-M2, ContraSoM incorporates modality-specific data augmentation techniques, including pixel-level transformations for images and diffusion-based feature-space augmentation specifically designed for the sparse and irregular structure of LiDAR point clouds. Our key contributions are summarized as follows:
@@
    We develop WiFo-M2, the first FM that can bring universal gains across diverse transceiver modules via multi-modal sensing. WiFo-M2 offers a new paradigm for sensing-aided communications field, avoiding the cumbersome steps traditionally required. As such, WiFo-M2 expands the set of transceiver modules that can be optimized through sensing, moving beyond the traditionally limited scope.
@@
    We propose a pre-training strategy based on contrastive learning, named ContraSoM. By pre-training on a large-scale multi-modal dataset, the ContraSoM strategy endows WiFo-M2 with the ability to mine universal OOB channel features during the inter-frame perception gap.
@@
    We apply WiFo-M2 to a wide range of existing transceiver design schemes. Experimental results demonstrate that the proposed WiFo-M2 can provide universal performance gains without requiring fine-tuning and exhibits strong cross-scenario generalization capabilities.
@@ II Related Work
II-A Sensing-Assisted Wireless Communications
@@
Multi-modal sensing has been increasingly utilized to improve wireless communication performance. Existing schemes are predominantly task-specific, requiring carefully customized data processing and network designs for each application. For example, the LE-CLN scheme [3] converts raw LiDAR data into the signal propagation feature superposed range map and customizes a feature extraction neural network (NN) to transform LiDAR into OOB features beneficial for channel estimation. Similarly, the H-MVMM scheme proposed in [8] is specifically designed for precoding, with carefully crafted data processing methods for Global Positioning System (GPS), LiDAR, and RGB images, as well as customized feature extraction and fusion NNs. The radar-based channel estimation method proposed in [4] extracts propagation delay, velocity, and angle parameters from radar data via clutter removal, 3D-FFT, and peak detection to guide sparse channel estimation. Such design principle leads to a fundamental limitation: the performance gain lacks universality and is only effective for a particular task, dataset, and transceiver setup.
@@
Another pressing issue is that current research remains concentrated on just a few transceiver modules whose outputs are strongly tied to user location, such as beam prediction and handover. To date, whether multi-modal sensing can benefit many other transceiver modules is still largely unexplored. The difficulty stems from the highly diverse and often latent relationships between sensing and transceiver modules. For example, temporal channel prediction (CP) relies on capturing the time-varying characteristics of scatterers; uplink–downlink CP concerns spatial consistency across frequency bands. Hence, crafting an dedicated sensing pipeline and mining task-specific features becomes labor-intensive. In this context, obtaining a universal beneficial feature from multi-modal sensing becomes an attractive topic. This study is the first to demonstrate that multi-modal sensing can boost channel interpolation and channel prediction via WiFo-M2, two core transceiver modules that had not been explored for such benefits until now.
II-B Contrastive Learning for Cross-Modal Alignment
@@
As a powerful self-supervised paradigm, CL has emerged as a foundational technique for aligning heterogeneous data modalities by learning a shared semantic space. Beyond vision-language domains, CL has been effectively adapted to align sensor modalities with complementary characteristics. Extending this idea, DiffCL [14] introduces a diffusion-based graph CL module that generates denoised visual- and text-aware views, encouraging the encoder to distil robust user–item representations even under severe interaction sparsity and multimodal noise. In medical image segmentation, CLMorph [15] integrates CL into a deformable registration framework to align an unaligned image with a reference image. By contrasting feature representations extracted from these images, it enhances the discriminative capacity of the network. In industrial time-series generation, MetaIndux-TS [16] employs a contrastive synthesis layer within its diffusion framework to enhance generation fidelity. This layer explicitly compares the synthesized sequences with the original noisy input, guiding the model to better capture fine-grained deviations and improve the overall quality of generated data.
@@
Building upon these advances, CL has also been introduced as an effective pre-training strategy in wireless communication community. For instance, CSI-CLIP proposed in [17] treats channel state information (CSI) and channel impulse response as naturally aligned pairs and performs cross-modal CL between them, enabling the model to capture generalizable channel features for improved positioning and beam management. LWLM [18] introduces Position-Invariant CL, which contrasts CSI samples collected from the same location under different base station (BS) configurations to learn robust and location-aware semantics specifically for localization tasks. Furthermore, in [19], a VLM-based beam prediction (BP) framework is proposed which utilizes CL to align image and LiDAR features, enforcing cross-modal consistency to enhance multi-modal fusion for beam selection. Most recently, the WMFM model [20] adopts a vanilla form of CL to align images with CSI, yet it still suffers from several limitations. Firstly, as in most prior work, it fails to resolve the intrinsic temporal mismatch between sensory data and CSI. Secondly, it utilizes solely the image modality and does not employ any data pre-processing and augmentation techniques, limiting its robustness and cross-domain generalization. Thirdly, the pre-training dataset covers only one scene and a single communication system configuration. Finally, it does not investigate how multi-modal sensing could benefit a wider range of transceiver modules and remains confined to Line of sight (LoS)/non-Line of sight (NLoS) classification and localization.
III System Model, Problem Description, and Challenges
III-A System Model and Problem Description
@@
We consider a typical intelligent system scenario in B5G/6G, e.g., an intelligent transportation system, where BS is equipped with multi‑modal sensors for environmental monitoring and situational awareness. In this work, we assume that multi‑modal sensing includes LiDAR point clouds and RGB images, denoted by $\bm{P}_{\text{raw}}\in\mathbb{R}^{N\times 3}$ and $\bm{I}\in\mathbb{R}^{3\times H\times W}$, respectively. We consider a multiple-input single-output orthogonal frequency division multiplexing (MISO-OFDM) communication system where BS is equipped with either a uniform linear array (ULA) or a uniform planar array (UPA), with the antenna size denoted by $N$ and the number of subcarriers by $K$. The BS serves $M$ single‑antenna users. Let $\mathbf{H}\in\mathbb{C}^{K\times N}$ represent CSI between the BS and a certain user.
In wireless communication systems, to establish a reliable link, the BS and users must go through a series of predefined steps, including but not limited to pilot transmission, channel estimation, full-band interpolation, precoder/decoder design, CP for high-mobility cases, and data detection. Each of these steps collectively constitutes the transceiver-design pipeline. The problem addressed in this paper is how to leverage multi‑modal sensing as external OOB information to assist various stages of the transceiver design, with the ultimate goal of streamlining and optimizing transceiver design to build a higher-performance communication system.
III-B Challenges
WiFo-M2 aims to serve as an FM capable of extracting universal OOB channel features from multi-modal sensing. Achieving this requires a self-supervised pre-training scheme that (i) learns from large-scale unlabeled data and (ii) aligns heterogeneous modalities. CL [13], which constructs representations by learning similarities between data samples, naturally meets the above requirements. CL [13] satisfies these criteria, yet its vanilla form is inadequate for WiFo-M2. Substantial modifications are needed in both the pre-training strategy and the model architecture to cope with the unique challenges posed by multi-modal sensing and wireless channels. The key obstacles are:
@@ III-B Challenges
(2) Temporal Misalignment Due to Heterogeneous Sampling Rates: The goal of sensing-aided communications is to provide OOB channel feature at the same measurement rate as the CSI estimation itself. However, a fundamental gap exists: CSI refreshes every few milliseconds, but cameras and LiDAR operate at low frame rates (e.g., 10-20 Hz). Most channel samples thus lack a concurrent sensing frame. Therefore, the model cannot rely on simply aligning sensory frames with coincident CSI samples. Instead, it has to effectively “fill in” the missing environmental context at the communication system’s operational tempo.
(3) Difficulty in Designing Effective Multi-modal Data Augmentation: Data augmentation is crucial for generating diverse, semantic-preserving views to learn robust representations. However, crafting augmentation strategies for complex, unstructured modalities like LiDAR point clouds is non-trivial. Simple operations like random point dropout or perturbation can easily corrupt geometric relationships, undermining both the authenticity of the simulation and the consistency required for CL. Effective augmentation method should be developed to maintain the data’s fundamental spatial and semantic structure to ensure the generated views remain valid positive samples.
Figure 2: Overall framework of WiFo-M2, illustrating the two‑stage pipeline: (left) pre-training using ContraSoM strategy, and (right) inference stage for optimizing transceiver design.
In this section, we introduce the core designs of WiFo‑M2: a unified sensing processing pipeline that addresses Challenge 1, the model architecture and feature extraction method that tackles Challenge 2, and a dedicated pre‑training strategy for WiFo‑M2 that equips WiFo‑M2 with essential capabilities while effectively resolving Challenge 3. The overall framework of WiFo‑M2 is depicted in Fig. 2.
IV-A Unified Sensing Processing Pipeline of WiFo-M2
IV-A1 Image Processing Pipeline
@@
Step 1: Initial Object Detection and Tracking. The raw image data is first subjected to a pixel-level augmentation pipeline (described in detail in Section IV-C1), and the resulting image is denoted as $\tilde{\bm{I}}$. The image processing pipeline begins with object detection to identify potential communication targets (e.g., vehicles in V2X communicatino scenarios) from $\tilde{\bm{I}}$. We assume the BS has prior knowledge of the receiver’s initial position through initial access procedures. Then, the advanced object detection model YOLOv8 [24] is used to detect all targets in the image and perform continuous tracking across frames.
@@
where $x_{\text{min}}=x_{c}-w/2$, $x_{\text{max}}=x_{c}+w/2$, $f_{x}$ is the focal length in pixels, and $c_{x}$ is the principal point offset along the x-axis.
@@
IV-A2 LiDAR Processing Pipeline

### v2

# WiFo-M2: Empower Wireless Communications With Plug-and-Play Environment Sensing via Foundation Model

@@ Abstract
The emerging convergence of next-generation wireless networks and agentic artificial intelligence (AI) is inspiring a new vision: embodied intelligent network entities utilize environmental sensing to refine their physical-layer (PHY) actions. Despite a growing body of preliminary work, prevailing small and task-specific AI models require extensive manual design of data pre-processing, network architecture, and fine-tuning, leaving them tightly coupled to particular PHY actions, system configurations, and deployment scenarios. To address this, we propose a paradigm shift with WiFo-M2, a foundation model that enables environment sensing to be easily integrated into PHY actions, delivering universal performance gains. To extract generalizable out-of-band (OOB) channel-aware features from environment sensing, we introduce ContraSoM, a contrastive pre-training strategy. Once pre-trained, WiFo-M2 infers future OOB channel-aware features from historical sensory data and strengthens feature robustness via modality-specific data augmentation. Experiments show that WiFo-M2 improves the performance of diverse PHY actions and demonstrates strong generalization to unseen scenarios.
@@ I Introduction
The 2030 era of networking will feature hundreds of billions of connected devices, immersive applications, and tight cyber physical loops that demand extreme reliability, ultra low latency, and sustainable efficiency [1]. Meeting these requirements calls for Agentic AI as a foundational paradigm in which autonomous networked agents perceive, reason, plan, and act [2]. Today, the application of Agentic AI in wireless networks is rapidly expanding, providing a robust framework for autonomous optimization and self-evolving in increasingly complex and dynamic environments [2]. Against this backdrop, Cheng et al. recently introduced Embodied Intelligent Wireless (EIW) [3] and envisioned it as the new paradigm for next-generation wireless communications. Under EIW, conventional network nodes such as base stations (BSs) and user equipment (UEs) evolve into intelligent network entities endowed with observation–decision–action capabilities.
@@
Driven by the rapid advancement of embodied intelligence, intelligent embodied network entities are increasingly equipped with comprehensive environmental sensing capabilities [4]. Motivated by this trend, the wireless communications community is actively exploring how these ubiquitous sensing sources can be leveraged to enhance physical layer (PHY) actions. Here, “actions” concretely refers to transceiver modules, e.g., channel estimation/prediction, and beamforming. To systematically give the design principle and underlying mechanisms of such sensing-refined PHY actions, Cheng et al. proposed the first framework of this field, Synesthesia of Machines (SoM) [4]. To date, numerous studies have investigated promising applications, such as using multi-modal sensing to boost channel estimation acquisition performance [7, 5, 6], to streamline beam selection and training [8, 9, 10, 11], and to optimize precoding [12].
However, as illustrated in Fig. 1, the field of sensing‑refined PHY actions remains dominated by traditional per‑task “small models” that deliver isolated gains but generalize poorly. They are scenario‑bound and configuration‑locked, lacking a cohesive mechanism to convert rich observations into consistently better PHY actions across environments. Moreover, the per-task paradigm keeps the research scope narrow and non-transferable. Each PHY action must be served by a customized data pipeline and model, driving up engineering cost and stalling large-scale deployment. Consequently, existing studies mostly address a few vision-friendly tasks such as beam prediction and handover, whereas broader needs like channel interpolation or prediction receive little attention. Finally, current BSs, though equipped with multi-modal sensors, still lack agentic‐AI capabilities. They passively fit sensing inputs to predefined outputs rather than learning latent scene dynamics and generally useful channel-aware features, leaving them unable to exploit observations under occlusion or to anticipate fine-grained out-of-band (OOB) features.
Figure 1: Comparisons between conventional RF-based/sensing-refined wireless communications paradigm, and the proposed FM-based approach of refining PHY actions with environment sensing under EIW paradigm.
@@
To propel current sensing-refined PHY research toward the EIW paradigm and the vision of Agentic AI, we advocate a foundation model (FM) based approach that equips intelligent entities with a universal, task-agnostic observation-to-action bridge. In this paper, we introduce WiFo-M2, the first universal FM that transforms multi-modal sensing into OOB channel-aware features, thereby serving as a plug-and-play performance booster across diverse PHY actions, as shown in Fig. 1. To establish one-to-one positive sample pairs between channel state information (CSI) and multi-modal sensory data, we first develop a unified multi-modal sensing processing pipeline. Then, we design the ContraSoM pre-training strategy, which is designed to align the high-dimensional representations extracted from multi-modal sensing with the corresponding CSI features in latent space. This alignment allows WiFo-M2 to extract universally beneficial channel-aware features for diverse PHY actions. To bridge the sampling rate gap among different modalities, WiFo-M2 is designed as a temporal sequence-to-sequence model. It ingests a fixed historical window of sensory data and is trained to extrapolate future channel-aligned features. This enables WiFo-M2 to provide channel-aware features even when fresh sensory data is temporarily unavailable, a practical issue largely overlooked in prior work. To enhance the robustness and representation learning capability of WiFo-M2, ContraSoM incorporates modality-specific data augmentation techniques, including pixel-level transformations for images and diffusion-based feature-space augmentation specifically designed for the sparse and irregular structure of LiDAR point clouds. Our key contributions are summarized as follows:
@@
    We develop WiFo-M2, the first FM that empowers multi-modal environmental sensing to enhance PHY actions in a scalable, plug-and-play manner. WiFo-M2 offers a scalable solution, avoiding the cumbersome, task-specific engineering traditionally required.
@@
    We propose a pre-training strategy based on contrastive learning (CL), named ContraSoM. By pre-training on a large-scale multi-modal dataset, the ContraSoM strategy endows WiFo-M2 with the ability to mine OOB channel-aware features during the inter-frame perception gap.
@@
    We validate the capabilities of WiFo-M2 across extensive PHY actions implemented via diverse schemes. Experiments show that WiFo-M2 empowers these PHY actions without requiring fine-tuning and exhibits strong cross-scenario generalization capabilities.
@@ II Related Work
II-A Sensing-Refined PHY Actions
@@
Multi-modal sensing has been increasingly utilized to improve PHY actions. Existing schemes are predominantly action-specific, requiring carefully customized data processing and network designs for each application. For example, the LE-CLN scheme [7] converts raw LiDAR data into the signal propagation feature superposed range map and customizes a feature extraction neural network (NN) to transform LiDAR into OOB features beneficial for channel estimation. Similarly, the H-MVMM scheme proposed in [12] is specifically designed for precoding, with carefully crafted data processing methods for Global Positioning System (GPS), LiDAR, and RGB images, as well as customized feature extraction and fusion NNs. The MSCP channel prediction framework proposed in [5] adopts a bespoke pipeline that first aligns camera images with pilots via angle transformation and path filtering, and then employs a task-specific cross-modal Transformer to regress channel parameters. Such design principle leads to a fundamental limitation: the performance gain lacks universality and is only effective for a particular PHY action, scenario, and transceiver setup.
@@
Another pressing issue is that current research remains concentrated on just a few PHY actions whose outputs are strongly tied to user location, such as beam prediction and handover. To date, whether environment sensing can benefit many other PHY actions is still largely unexplored. The difficulty stems from the highly diverse and often latent relationships between sensing and PHY actions. For example, temporal channel prediction (CP) relies on capturing time-varying characteristics of scatterers; uplink–downlink CP concerns spatial consistency across frequency bands. Hence, crafting a dedicated sensing pipeline and mining task-specific features becomes labor-intensive. In this context, how to obtain a universal OOB feature from multi-modal sensing and use it to boost a wide range of PHY actions becomes a critical challenge.
II-B CL-Based Foundation Model for Cross-Modal Alignment
@@
WiFo-M2 is envisioned to extract universal OOB channel-aware features from multi-modal environment sensing. Achieving this requires a self-supervised pre-training scheme that (i) learns from large-scale unlabeled data and (ii) aligns heterogeneous modalities. CL [13], which constructs representations by learning similarities between data samples, naturally meets the above requirements. Recently, beyond vision-language domains, CL has been effectively adapted to align sensor modalities with complementary characteristics. DiffCL [14] introduces a diffusion-based graph CL module that generates denoised visual- and text-aware views, encouraging the encoder to distill robust user–item representations even under severe interaction sparsity and multi-modal noise. In medical image segmentation, CLMorph [15] integrates CL into a deformable registration framework to align an unaligned image with a reference image. By contrasting feature representations extracted from these images, it enhances the discriminative capacity of the network. Similarly, in multi-media recommendation, graph-based contrastive methods [16] are designed to disentangle and align user preferences across visual, acoustic, and textual modalities.
@@
Building upon these advances, CL has also been introduced as an effective pre-training strategy in wireless communication community. CSI-CLIP proposed in [17] treats CSI and channel impulse response as naturally aligned pairs and performs cross-modal CL between them, enabling the model to capture generalizable channel features for improved positioning and beam management. LWLM [18] introduces position-invariant CL, which contrasts CSI samples collected from the same location under different BS configurations to learn robust and location-aware semantics specifically for localization tasks. Furthermore, in [19], a VLM-based beam prediction (BP) framework is proposed which utilizes CL to align image and LiDAR features, enforcing cross-modal consistency to enhance multi-modal fusion for beam selection. Most recently, the WMFM model [20] adopts a vanilla form of CL to align images with CSI, yet it still suffers from several limitations. Firstly, as in most prior work, it fails to resolve the intrinsic temporal mismatch between sensory data and CSI. Secondly, it utilizes solely the image modality and does not employ any data pre-processing and augmentation techniques, limiting its robustness and cross-domain generalization. Thirdly, the pre-training dataset covers only one scene and a single communication system configuration. Finally, it does not investigate how multi-modal sensing could benefit a wider range of PHY actions and remains confined to Line of sight (LoS)/non-Line of sight (NLoS) classification and localization.
III System Model
As illustrated in Fig. 2, we consider an intelligent transportation scenario, where the BS is an intelligent agent that serves $M$ single‑antenna users and senses the physical world through multi‑modal sensors. In this work, we assume that multi‑modal environment sensing includes LiDAR point clouds and RGB images, denoted by $\bm{P}_{\text{raw}}\in\mathbb{R}^{N\times 3}$ and $\bm{I}\in\mathbb{R}^{3\times H\times W}$, respectively. We assume the camera operates at $40$ fps, while the LiDAR scans at $20$ Hz.
@@ III System Model
We consider a multiple-input single-output orthogonal frequency division multiplexing (MISO-OFDM) communication system under EIW paradigm where BS is equipped with either a uniform linear array (ULA) or a uniform planar array (UPA), with the antenna size denoted by $N$ and the number of subcarriers by $K$. Let $\mathbf{H}\in\mathbb{C}^{K\times N}$ represent CSI between the BS and a certain user.
Figure 2: An overview of the considered scenario.
Figure 3: Overall framework of WiFo-M2, illustrating the two‑stage pipeline: (left) pre-training stage with ContraSoM strategy realizing SoM processing, and (right) inference stage for optimizing various PHY actions in a plug-and-play manner.
In this section, we first analyze the fundamental challenges faced by developing WiFo-M2 to achieve scalable environment sensing-refined PHY actions for intelligent network entities. Then, we introduce the core designs: the unified environment sensing processing pipeline, the architecture and feature extraction method of the WiFo-M2, and the ContraSoM pre‑training strategy. The overall framework of WiFo‑M2 is depicted in Fig. 3.
IV-A Challenges
To realize the goal of scalable sensing-refined PHY actions, the primary obstacles arise from three key aspects: environment sensing data processing, foundation model design, and the pre-training strategy. These are detailed as follows:
@@ IV-A Challenges
(2) Temporal Misalignment Due to Heterogeneous Sampling Rates: Wireless channels vary every few milliseconds, whereas intelligent network entities can only observe the surrounding environment at low frame rates (e.g., 1010-2020 Hz). That is, most PHY actions lack concurrent environmental observations during execution. Therefore, WiFo-M2 cannot simply rely on aligning sensory frames with coincident CSI samples. Instead, it has to effectively “fill in” the missing environmental context at the communication system’s operational tempo.
(3) Effective Multi-modal Data Augmentation Technique: Data augmentation is crucial for generating diverse, semantic-preserving views to learn robust representations. However, crafting augmentation strategies for complex, unstructured modalities like LiDAR point clouds is non-trivial. Simple point dropout or perturbation can easily corrupt geometric relationships, undermining the consistency required for CL. Effective augmentation methods should be developed to maintain the data’s fundamental spatial and semantic structure to ensure the generated views remain valid positive samples.
IV-B Unified Environment Sensing Processing Pipeline
IV-B1 Image Processing Pipeline
@@
Step 1: Initial Object Detection and Tracking. The raw image data is first subjected to a pixel-level augmentation pipeline (described in detail in Section IV-D1), and the resulting image is denoted as $\tilde{\bm{I}}$. The image processing pipeline begins with object detection to identify potential communication targets (e.g., vehicles in V2X communication scenarios) from $\tilde{\bm{I}}$. We assume the BS has prior knowledge of the receiver’s initial position through initial access procedures. Then, the advanced object detection model YOLOv8 is used to detect all targets in the image and perform continuous tracking across frames.
@@
where $x_{\text{min}}=x_{c}-w/2$, $x_{\text{max}}=x_{c}+w/2$, $f_{x}$ is the focal length in pixels, and $c_{x}$ is principal point offset along x-axis.
@@
IV-B2 LiDAR Processing Pipeline